solve for x log 5x - log 5 3=2
I got log 5 x/3=2
x=2/3
is this right?

Answer by drk(1908)

solve for x:
log_5(x) - log_5(3)=2
Step 1: notice we are subtracting log's. The rule is if you subtract logs, then divide the information as
log_5(x/3) = 2.
step 2: rewrite as an exponent as:
5^2 = x/3
Step 3: solve for x:
x = 75.